OpenStudy (anonymous):

Solve for R:

OpenStudy (anonymous):

\[V= 4/3 \pi r ^{3}\]

OpenStudy (anonymous):

@shae.jones

OpenStudy (anonymous):

A. \[3V \pi / 2\]

OpenStudy (anonymous):

B. \[( \sqrt[3]{3V/4\pi})\]

OpenStudy (anonymous):

C. \[\left( \sqrt{3V \pi/4} \right)\]

OpenStudy (anonymous):

wat does V equal

OpenStudy (anonymous):

D. \[r = 3V \pi / 4\]

OpenStudy (anonymous):

* A, B, & C start with r = *

OpenStudy (anonymous):

@shae.jones : no it doesnt have to = anything .

OpenStudy (anonymous):

@FlufflePuffle

OpenStudy (anonymous):

so a,b,and c are answers choices

OpenStudy (anonymous):

yes . A,B,C, & D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my answer would be D